OpenStudy (anonymous):

I have a function f(x) = sqrt(5x+1) and I'm trying to find f'(a) using the definition: lim h->0 f(a+h) - f(a)/h I'm having a hard time simplifying the expression. Help!

OpenStudy (anonymous):

Okay. You have: \[\lim_{h \rightarrow 0}\frac{\sqrt{5x+5h+1}-\sqrt{5x+1}}{h}\] Now let me look at it :P

OpenStudy (anonymous):

I got it :P I think. Multiply the top and bottom by: \[\sqrt{5x+5x+1}+\sqrt{5x+1}\] After that, see if you can figure it out :P

OpenStudy (anonymous):

lim┬(h→0)⁡〖([√(5(a+h)+1)]- [√(5a+1)])/h〗

myininaya (myininaya):

rationalize!

myininaya (myininaya):

ratiaonlize the numerator!

myininaya (myininaya):

multiply by our favorite thingy the conjugate of the top

myininaya (myininaya):

if we do it to the top we have to do it the bottom

OpenStudy (anonymous):

malevolence has it . you have to do some algebra, but the numerator will be \[5x+5h+1-(5x+1)=5h\]

myininaya (myininaya):

the h will cancel because u have one on bottom too

OpenStudy (anonymous):

Then you get: \[\lim_{h \rightarrow 0}\frac{5x+5h+1-5x-1}{h(\sqrt{5x+5h+1}+\sqrt{5x+1})}\] Everything cancels in the top leaving a 5h. Divide out the h in the top and bottom. Then you take the limit and the denominator becomes: \[\sqrt{5x+1}+\sqrt{5x+1}=2\sqrt{5x+1}\] Giving! \[\frac{5}{2\sqrt{5x+1}}\]

OpenStudy (anonymous):

and the denominator will be \[h(\sqrt{5x+5h+1}+\sqrt{5x+1})\]

myininaya (myininaya):

appletastic!

OpenStudy (anonymous):

cancel the h wow! nice latex malevolence!

OpenStudy (anonymous):

I was trying to type quickly :P Got you and myininaya jumping down my throat xPPP

myininaya (myininaya):

lol

OpenStudy (anonymous):

\[\color{green}{\text{appletini}}\]

myininaya (myininaya):

omg satellite you will not believe this but malevolence corrected me earlier

OpenStudy (anonymous):

latex color race this fall

myininaya (myininaya):

i tried to tell him it was just a typo like you always say

OpenStudy (anonymous):

\[\color{blue}{\text{you probably needed it}}\]

OpenStudy (anonymous):

\[\huge\int\limits x^2dx\]

OpenStudy (anonymous):

when i do it: typo when you do it: error

myininaya (myininaya):

wow thats big

OpenStudy (anonymous):

LARGE!

myininaya (myininaya):

the satellite signal is out! come on i will be your sidekick

OpenStudy (anonymous):

\[\huge\int\limits \int\limits_D \int\limits \frac{P(z)}{Q(z)}dV\]

OpenStudy (anonymous):

\[\huge\text{mine's bigger}\]

OpenStudy (anonymous):

\[\huge\color{red} {redder}\]